40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 326 of 349  Not logged in ELOG logo
ID Date Author Type Category Subject
  1208   Tue Dec 30 18:51:18 2008 rana,yoichiConfigurationElectronicsIlluminator Power Supply reset
We noticed that none of the illuminators were working.

The switches were off on all the ports. After turning them on it still didn't work.

The +24 V Sorensen power supply which powers all of the illuminators had its OVP light on.
We turned it off, ramped the voltage to zero, turned it back on, and then went back to +24 V.

We then checked the operation of the illuminators; ETMY is still MIA.

Each of the illuminators sucks ~0.6-0.7 A when the (unlabeled) rheostat knob panel is set
to the "25" setting.

It seems pretty unwise, in the EMI sense, to be sending Amps of unshielded, high current,
switching supply outputs for 40m down the arms. This creates a huge antenna for radiating
the switching noise. I hereby assign minus 5 points to whoever designed this system.

Illuminator Upgrade:
- Use LEDs of a wavelength that the OSEMs don't see. LEDs are also cool so that the
  Suspension won't drift in alignment.

- Use 2 power supplies so that the power is balanced.

- Make is +/-12 V twisted AWG 14 wire so that the EMI is contained. Should also
  be shielded cable.
  1207   Mon Dec 29 21:51:02 2008 YoichiConfigurationComputersWeb server on nodus
The apache on nodus has been solely serving for the svn web access.
I changed the configuration and all files under /cvs/cds/caltech/users/public_html/ can be seen under
https://nodus.ligo.caltech.edu:30889/

The page is not password protected, but you can add a protection by putting an appropriate .htaccess
in your directory.
For the standard LVC password, put the following in your .htaccess
AuthType Basic  
AuthName "LVC password"
AuthUserFile /cvs/cds/caltech/apache/etc/LVC.auth
Require valid-user
  1206   Mon Dec 29 21:38:57 2008 YoichiUpdateComputersSnapshots of MEDM screens
I wrote scripts to take snapshots of MEDM screens in the background.
These scripts work even on a computer without a physical display attached.
You don't need to have X running.
So now the scripts run on nodus every 5 minutes from cron.
The screen shots are saved in /cvs/cds/caltech/statScreen/images/

There is a wiki page for the scripts.
http://lhocds.ligo-wa.caltech.edu:8000/40m/captureScreen.sh

Someone has to make a nice web page summarizing the captured images.
  1205   Mon Dec 29 18:01:07 2008 AidanUpdateAuxiliary lockingUpdated 40m Upgrade Document T080074-00-R

Added a paragraph to the 40m Upgrade document describing the fiber stabilization and frequency doubling proposed for auxiliary locking.

Also added a complete diagram of the fiber stabilization and a draft sketch of the frequency doubling.

Uploaded to https://nodus.ligo.caltech.edu:30889/svn/trunk/docs/upgrade08/ via svn.
  1204   Wed Dec 24 12:46:54 2008 YoichiUpdateComputersTest points are back
Rob told me how to restart the test point manager.
It runs on fb40m and actually there is an instruction on how to do that in the Wiki.
http://lhocds.ligo-wa.caltech.edu:8000/40m/Computer_Restart_Procedures#fb40m

I couldn't find the page because when I put a keyword in the search box on the upper right
corner of the Wiki page and hit "enter", it only searches for titles. To do a full text
search, you have to click on the "Text" button.

Anyway, now the test points are back.
  1203   Wed Dec 24 10:33:24 2008 YoichiUpdateComputersSeveral fixes. Test point problem remains.
Yesterday, I fixed several remaining problems from the power failure.

I found a LEMO cable connecting the timing board to the Penteks was lose on the c1susvme1 crate.
After I pushed it in, the DMA error has not occured on c1susvme1.

I logged into op340m using a Null Modem Cable.
The computer was failing to boot because there were un-recoverable disk errors by the automatic fsck.
I run fsck manually and corrected some errors. After that, op340m booted normally and now it is working fine.
Here is the serial communication parameters I used to communicate with op340m:
>kermit      (I used kermit command for serial communication.)
>set modem type none
>set line /dev/ttyS0     (ttyS0 should be the device name of your serial port)
>set speed 9600
>set parity none
>set stop-bits 1
>set flow-control none
>connect

After fixing op340m, the MC locked.
Then I reset the HV amps. for the steering PZTs.
Somehow, the PZT1 PIT did not work. But after moving the slider back and forth several times, it started to work.

I reset the mechanical shutters around the lab.

I went ahead to align the mirrors. The X-arm locked but the alignment script did not improve
the arm power.
I found that test points are not available. (diag said test point management not available).
Looks like test point manager is not running. Called Rolf, but could not reach him.
I'm not even sure on which machine, the tp manager is supposed to be run.
Is it c0daqawg ?
  1202   Tue Dec 23 10:35:40 2008 YoichiUpdateComputersRFM network breakdown mostly fixed
Rana, Rolf, Alberto, Yoichi

The source of the problem was the RFM bypass box, as expected.
Rana pointed out that the long cable I used to bring the 5V from the Sorensen to the box
may cause a large voltage drop considering that the box is sucking ~3A.
So we connected the cable to another power supply (5V/5A linear power supply).
Then the LEDs on the bypass box turned green from red, and everything started to work.

A weired thing is that when I connected the cable to the wrong terminals of the power supply which
have lower current supply capabilities, the supply voltage dropped to 3V, but still the LEDs on the bypass box
turned green. This means the bypass box can live with 3V.
I noticed that there is a long cable from the Sorensen to the cross connect on the side of the rack, where I
connected my cable to the bypass box. This long cable had somewhat large resistance (1 or 2 Ohms) and dropped
the supply voltage to less than 3V ?
Anyway, the bypass box is now on a temporary power supply. Alberto was assigned a task to find a replacement power
supply.

There are two remaining problems.
c1susvme1 fails to start often claiming a DMA error on a Pentek. After several attempts, you can start the machine,
but after a while (1 hour ?) it fails again.
op340m is not responding to ssh login. It responds to ping.
We hooked up a monitor and keyboard (USB because the machine does not have a PS/2 port) to it and rebooted.
At the boot, it briefly displays a message "No keyboard, try TTYa", but after that no display signal.
Steve found me a serial cable. I will try to login to the machine using the serial port.

  1201   Mon Dec 22 13:48:22 2008 YoichiUpdateComputersRFM network bypass box's power supply is dead
As a temporary fix, I cut the cable of the power supply and connected it to the Sorensen power supply +5V on the rack.
Now, the RFM bypass box is powered up, but some LEDs are red, which looks like a bad sign.
I restarted all the FE computers, but this time I got errors during the execution of the startup commands in the VxWorks machines.
The errors are "General Protection Fault" or "Invalid Opcode".
The linux machines do not show errors but still the status lights in EPICS are red.
We need Alex's help. He did not answer the phone, so Alberto left a voice mail.
  1200   Sun Dec 21 14:18:04 2008 YoichiUpdateComputersRFM network bypass box's power supply is dead
I restarted the front-end computers by power cycling them one-by-one.
After issuing startup commands, most of them started normally at least by looking
at the output from telnet/ssh.
However, the status monitors of the FE computers on the EPICS screen are still red.
I noticed that all the LEDs on the VMIC 5594 RFM network bypass box are off.
According to the labels, fb40m, c0daqctrl, c0dcu are connected to the box.
This means (I believe) c1dcuepics cannot access the RFM network. So we have no control over
the FE computers through EPICS.

I pushed the reset button on the box, power cycled it, but nothing changed.
I checked the fuse and it was OK. Then I found that the power supply was dead.
It is a small AC adapter supplying +5VDC with a 5-pin DIN like connector.
We have to find a replacement.
  1199   Sun Dec 21 13:00:06 2008 steveUpdateall down cond.vac and laser back on
There was a power outage sometimes early Saturday.

All things are down.

The Maglev was started and reached normal operating condition.
V1 was opened at P1 3.8 mTorr and cc1 is back to 3e-6 Torr now

The MOPA was turned on.
Ion pump HV on to FSS cavity.
  1198   Sat Dec 20 23:37:43 2008 robOmnistructureGeneralSaturday Night Fever after presumed power failure

Just came by to pick something up...

... alarm handlers screeching...

... TP1 failure--closing V1... call Steve... Steve says ok till tomorrow...

... all front ends down (red)...

... all suspensions watchdogged...

... all (I think) servos off...

... PSL shutter closed ...

... chiller at 15C ... I turned it off to prevent condensation in PA...

... MOPA shutter closed... turned off key on Lightwave power supply

... good luck all, and happy holidays!
  1197   Fri Dec 19 16:38:09 2008 steveUpdateLSCall optlevs centered
All optlevs were centered after full alignment.

Qpd sums are:
ETMX 12,229 counts
ITMX 9,932
ETMY 12,043
ITMY 4,362
BS 1,880
PRM 1,423
SRM 11,641
  1196   Fri Dec 19 14:35:58 2008 Yoichi AlbertoUpdateIOOMC WFS and IOO-POS QPD re-centering
For the past two days, the MC alignment has kept drifting.
This morning, the MC alignment was so bad that it wouldn't lock to the TEM00 mode.
We aligned the MC mirrors manually until the reflection looks like a nice bull's-eye (the WFSs were off at this moment).
Then we un-locked the MC and centered the beams on the WFS QPDs.
Since the QPDs were saturated with the full laser power falling on them, I reduced the PSL power by turning the HWP after the MOPA.
After this, we turned on the WFSs and everything looks normal now.
We will see the trend of the MC related channels to monitor the drift.

Although unlikely, it might be caused by the drift of the input beam to the MC.
We found that the IOO-POS QPD was mis-centered and saturating.
We replaced the BS picking up the beam for the QPD from 33% reflection to 10% one. The QPD was still saturated.
So we put the 33% BS in the beam path to the QPD to further reduce the power. The beam kicked by the 33% BS
is dumped to a black aluminum plate. We should use a better beam dump later.
Now the IOO-POS QPD should tell us some information about the beam pointing of the PSL, though it has no sensitivity
to the relative motion of the PSL table to the vacuum chambers.
  1195   Fri Dec 19 11:29:16 2008 Alberto, YoichiConfigurationMZMZ Trans PD
Lately, it seems that the matching of the input beam to the Mode Cleaner has changed. Also, it is drifting such that it has become necessary to continuously adjust the MC cavity alignment for it to lock properly.

Looking for causes we stopped on the Mach Zehnder. We found that the monitor channel:
C1:PSL-MZ_MZTRANSPD

which supposedly reads the voltage from some photodiode measuring the transmitted power from the Mach Zehnder, is totally unreliable and actually not related to any beam at all.

Blocking either the MZ input or output beam does not change the channel's readout. The reflection channel readout responds well, so it seems ok.
  1194   Fri Dec 19 11:18:52 2008 AlbertoConfigurationGeneralMode Cleaner Temperature Monitor
I reduced from 10 to 5 the gain of the SR560 that Caryn has set up after the lock-in amplifier nest to the PSL rack because the overload LED was flashing.
  1193   Thu Dec 18 19:15:54 2008 Alberto, YoichiConfigurationSUSMode Cleaner Cavity Alignment

Quote:
This morning I found the MC locked to the 10 mode. When I locked it on the 00 mode, it was unstable and eventually it always got locked to the wrong mode.

I looked at the Drift Mon MEDM screen, which shows a reference record for position, pitch and yaw of each mirror, and I found that the MC optics were in a different status. Moving the sliders of the mirrors' actuators, I brought them back to the reference position. Then the lock got engaged and it was stable, although the MC reflection from the photodiode, with the wave front sensors (WFS) off, was about 2V. That's higher than the 0.5V the it could get when we aligned the cavity and the input periscope last time.

With the WFS on, the reflection dropped to 0.3V and, so far, the the cavity has been stably locked.


This evening the mode cleaner was again locking on a higher mode so we tweaked the mirrors' actuators by their sliders on the MEDM screen until we improved the reflection to 0.3V.

Then we went inside and, on the AS table, we centered the beam on the wave front sensors.

Now the mode cleaner is locked, the reflection is less than 0.3V and the transmission about 3V, tha is it is in ideal conditions. We'll see if it holds.
  1192   Thu Dec 18 12:52:00 2008 AlbertoConfigurationSUSMode Cleaner Cavity Alignment
This morning I found the MC locked to the 10 mode. When I locked it on the 00 mode, it was unstable and eventually it always got locked to the wrong mode.

I looked at the Drift Mon MEDM screen, which shows a reference record for position, pitch and yaw of each mirror, and I found that the MC optics were in a different status. Moving the sliders of the mirrors' actuators, I brought them back to the reference position. Then the lock got engaged and it was stable, although the MC reflection from the photodiode, with the wave front sensors (WFS) off, was about 2V. That's higher than the 0.5V the it could get when we aligned the cavity and the input periscope last time.

With the WFS on, the reflection dropped to 0.3V and, so far, the the cavity has been stably locked.
  1191   Tue Dec 16 19:06:01 2008 YoichiUpdatePSLReference cavity ring down repeated many times
Today, I repeated the reference cavity ring down measurement many times to see how much the results vary.

I repeated the ring down for 20 times and the first attachment shows the comparison of the measured and estimated cavity transmission power.
The blue curve is the measured one, and the red curve is the estimated one. There are only 10 plots because I made a mistake when transferring data
from the oscilloscope to the PC, and one measurement data was lost.

The second attachment shows the histogram of the histogram of the estimated cavity pole frequencies.
I admit that there are not enough samples to treat it statistically.
Anyway, the mean and the standard deviation of the estimated frequencies are 47.6kHz and 2.4kHz.
Assuming a Gaussian distribution and zero systematic error, both of which are bold assumptions though, the result is 47.6(+/-0.6)kHz.

Now I removed the Pockels Cell from the RC input beam path.
I maximized the transmission by tweaking the steering mirrors and rotating the HWP.
Since the transmission PD was saturated without an ND filter on it, I reduced the VCO RF power slider to 2.85.
Accordingly, I changed the nominal common gain of the FSS servo to 10.5dB.
Attachment 1: RC_Ringdown_Estimates.png
RC_Ringdown_Estimates.png
Attachment 2: Cavity_Pole_Histogram.png
Cavity_Pole_Histogram.png
  1190   Fri Dec 12 22:51:23 2008 YoichiUpdatePSLReference cavity ring down measurement again
Bob made new HV-cables with HV compatible coaxes. The coax cable is rated for 2kV, which was as high as Bob
could found. I used it with 3kV hoping it was ok.
I also put a series resistor to the pockels cell to tame down the ripples I saw in elog:1136.

Despite those efforts, I still observed large ringings.
I tried several resistor values (2.5k, 1k, 330ohm), and found that 330ohm gives a slightly better result.
(When the resistance is larger, the edge of the PBS Refl. becomes dull).
Since the shape of the ringing does not change at all even when the pulse voltage is lowered to less than 1kV,
I'm now suspicious of the DEI pulser.

Anyway, I estimated the cavity pole using the MATLAB's system identification toolbox again.
This time, I locked the reference cavity using only the PZT feedback, which makes the UGF about a few kHz.
So, within the time scale shown in the plot below, the servo does not have enough time to respond, thus the laser
frequency stays tuned with the cavity. This was necessary to avoid non-linear behavior of the transmitted power
caused by the servo disturbing the laser frequency. With this treatment, I was able to approximate the response of
the cavity with a simple linear model (one pole low-pass filter).

MATLAB estimated the cavity pole to be 47.5kHz.
The blue curve in the plot is the measured RC transmitted power.
The incident power to the cavity can be inferred from the inverse of the red curve (the PBS reflection power).
The brown curve is the response of the first order low-pass filter with fc=47.5kHz to the input power variation.
The blue and brown curves match well for the first 10usec. Even after that the phases match well.
So the estimated 47.5kHz is probably a reasonable number. I don't know yet how to estimate the error of this measurement.

According to http://www.ligo.caltech.edu/~ajw/PSLFRC.png the designed transmission of the reference cavity mirrors is 300ppm (i.e.
the round trip loss (RTL) is 600ppm).
This number yields fc=35kHz. In the same picture, it was stated that fc=38.74kHz (I guess this is a measured number at some point).
The current fc=47.5kHz means, the RTL has increased by 200ppm from the design and 150ppm from the time fc=38.74kHz was measured.
Attachment 1: RC-Ringdown.png
RC-Ringdown.png
  1189   Tue Dec 9 10:48:17 2008 CarynSummaryGeneralcalibrating the jenne laser: impedance mismatch?

We sent RFout of network analyzer to a splitter, with one side going back to the network analyzer and the other to the laser modulation input. We observed a rippled transfer function through the splitter. The ripple is probably due to reflection due to an impedance mismatch in the laser.
Attachment 1: reflection.png
reflection.png
  1188   Mon Dec 8 17:50:21 2008 YoichiUpdateSUSITMY drift
The suspension drift monitor shows that the ITMY alignment was shifted after the earthquake.
Looks like only the UL sensor had a step at the earthquake (see the attachment 1).
So it is probably an electronics problem.
I pushed in the cable between the rack and the ITMY satellite amplifier, but no change observed.
Actually, the ITMY-UL sensor looks like it has been dead before the earthquake.
The second attachment shows a long-term trend of the UL sensor.
The sensor output had been around zero since Nov. 17th.
When I disabled the output of the UL sensor, the sus-drift-mon fields turned green.
So I think the drift-mon's reference values are wrong, and currently the ITMY is in a good alignment.

I also attached the free-swing measurements of the ITMY taken on Aug. 18th and today.
There is no notable change in the resonant frequencies.
Attachment 1: ITMY-OSEMs.png
ITMY-OSEMs.png
Attachment 2: ITMY-UL.png
ITMY-UL.png
Attachment 3: ITMY-08-18.pdf
ITMY-08-18.pdf
Attachment 4: ITMY-12-08.pdf
ITMY-12-08.pdf
  1187   Mon Dec 8 11:54:27 2008 YoichiUpdateGeneralIFO mirrors aligned
This morning, I re-aligned the IFO mirrors to see if they were badly moved by the earthquake.
The both arms locked just by the restoring scripts, but the transmission was about 0.7. So I aligned them
with the dithering scripts.
To lock the PRMI, I had to manually tweak the PRM alignment. After running the dithering script, the SPOB
went up to 1200.
I also had to tweak the SRM to get the DRMI locked. After the dithering script, the SPOB was 4200 and REFL166Q
was 3000.
  1186   Mon Dec 8 11:41:27 2008 YoichiSummaryVACThe rough pump for the TP2 replaced
Bob, Yoichi

The foreline pressure of the TP2 (the foreline pump for the main mag-lev turbo (TP1)) was at 2.8torr this morning
when Bob came in.
Looked like the foreline pump (Varian SH-110) was leaking.
Bob started the backup rough pump in parallel with the "leaking" one to keep the foreline pressure low.
We then closed the valve 4 (between TP2 and TP1) and stopped the TP2 and the SH-110.
We replaced SH-110 with another one, but still the foreline pressure was high.
So we replaced it with yet another one. We also changed the quick coupling fasteners on the SH-110 and wiped the O-rings.
This time, it worked fine and the foreline pressure dropped to around 38 mTorr.

Since there is no valve between the TP2 and the SH-110, we could not keep the TP2 running while we were replacing the
problematic SH-110. This means the TP1 was running without a foreline pump during the work. We tried to minimize the
down time of the TP2. The temperature of the TP1 was 33.6C before we stopped the TP2 and it went up to 37.3C during the
work. It is now coming down to the original temperature.

Since we don't know if the problem was caused by bad SH-110s or leaking quick couplings, Bob is checking these apparently
"leaking" SH-110s.
  1185   Mon Dec 8 00:10:42 2008 carynSummaryGeneralcalibrating the jenne laser
I apologize in advance for the long list of numbers in the attachment. I can't seem to make them hide for some reason.

So, since Jenne's laser will probably be used for the Stoch mon calibration, Alberto and I took some measurements to calibrate Jenne's laser.
We focused the beam onto the New Focus RF 1GHz photodetector that we stole from rana's lab (powered with NewFocus power 0901). Measured the DC output of the photodetector with scope. Aligned the beam so DC went up (also tried modulating laser at 33MHz and aligning so 33MHz peak went up). Hooked up the 4395a Spectrum/Network Analyzer to the laser and to the AC out of the photodetector (after calibrating Network analyzer with the cables) so that the frequency response of the laser*photodetector could be measured.
(Note: for a while, we were using a splitter, but for the measurements here, I got rid of the splitter and just sent the RFout through the cables to channel A for the calibration, sent RFout to the laser and photodetector to channel A for the measurement)

Measured the frequency response. At first, we got this weird thing with a dip around 290MHz (see jcal_dip_2_norm.png below).
After much fiddling, it appeared that the dip was from the laser itself. And if you pull up just right on the corner of this little metal flap on the laser (see picture), then the dip in the frequency response seems to go away and the frequency response is pretty flat(see jcal_flat_3_norm below). If you press down on the flap, the dip returns. This at least happened a couple of times.
Note that despite dividing the magnitude by the DC, the frequency responses don't all line up. I'm not sure why. In some cases the DC was drifting a bit(I presume the laser was coming out of alignment or decided to align itself better) and maybe with avgfactor=16, and measuring mean DC on the scope, it made the DC meas not match up the the frequ resp meas...
I've attached the data for the measurements made (I'm so sorry for all the #'s. I can't figure out how to hide them)
name/lasercurrent/DC/analyzer SourcePower/analyzer avgfactor
jcal7_1/I=31.7mA/DC=-4.41/SourcePower=0dBm/avgfactor=16
jcal7_2/I=31.7mA/DC=-1.56/SourcePower=0dBm/avgfactor=none
jcal8_1/I=31.7mA/DC=-4.58/SourcePower=0dBm/avgfactor=16
jcal8_2/I=31.7mA/DC=-2.02/SourcePower=0dBm/avgfactor=16
jcal8_3/I=31.7mA/DC=-3.37/SourcePower=0dBm/avgfactor=16
Note also that the data from the 4395a seems to have column1-frequency, column2-real part, column3-imaginary part...I think. So, to calculate the magnitude, I just took (column2)^2+(column3)^2.


To get sort of an upper-bound on the DC, I measured how DCmax varied with laser current, where DCmax is the DC for the best alignment I could get. After setting the current, the laser was modulated at 33MHz and the beam was aligned such that the 33MHz peak in the photodetector output was as tall as I could manage. Then DC was measured. See IvsDCmax.png. Note the DC is negative. I don't know why.

Also, the TV's don't look normal, the alarm's going off and I don't think the mode cleaner's locked.
Attachment 1: IvsDCmax.png
IvsDCmax.png
Attachment 2: data.tar.gz
Attachment 3: jcal_dip_2_norm_log.png
jcal_dip_2_norm_log.png
Attachment 4: jcal_flat_3_norm_log.png
jcal_flat_3_norm_log.png
  1184   Sun Dec 7 16:12:53 2008 ranaUpdateDAQbooted awg
because it was red
  1183   Sun Dec 7 16:02:46 2008 ranaUpdateGeneralMag 5.1 EQ halfway to Vegas
There was a Mag 5.1 EQ Friday night at 8:18 PM.

It tripped most of our optics. They all damped down passively except for MC2. Further more, ITMY seems to have come back to a different place.

Don't know why MC2 was so upset but I think maybe its watchdog didn't work correctly and the side loop is unstable when there are
large motions. After I lowered the side gain by 10x and waited a few minutes it came back OK and the MC locked fine.

I have just now put all the WDs into the Shutdown state so that we can collect some hours of free swinging data to see if there's been
any damage. Feel free to redamp the optics whenever you need them. Someone should do the eigenfrequency check in the morning and compare
with our table of frequencies in the wiki.

I excited the optics using the standard SUS/freeswing-all.csh script. Here's the output:
Excited all optics
Sun Dec  7 16:07:32 PST 2008
Attachment 1: g.png
g.png
Attachment 2: Untitled.png
Untitled.png
  1182   Fri Dec 5 21:31:11 2008 YoichiUpdateIOOdrum modes observable without excitation
The calibration of the MC_F feedback is posted in elog:1032.
I'm not sure where Caryn took MC signal, but if you take the signal from the servo out BNC on the MC board, it
directly corresponds to the voltage sent to the FSS VCO.
The DC calibration of the VCO is 1.75MHz/V. Since the AOM is the double-pass, the PSL frequency
change is 3.5MHz/V. At frequencies above 40Hz, the VCO calibration drops by a factor of 39/1000,
because of the pole/zero at 1.6Hz/40Hz in the VCO box.
So at the frequencies of interest (around 30kHz), the servo out voltage can be converted to the PSL frequency
change by 0.137MHz/V.
Since 30kHz is still within the bandwidth of the MC servo, the feedback signal should correspond to the actual
length change of the MC. So the above calibration factor can be used to calibrate Caryn's measurement and check
what Rana suggested.
  1181   Fri Dec 5 20:40:38 2008 YoichiHowToComputersElog multi-keyword search
The current Elog search allows you to look for only one keyword in the text.
You cannot search for two keywords by simply separating them with a white space.
That is, a search term "abc def" matches a literal "abc def", not a text containing "abc" and "def".
This is extremely annoying. However, there are still some ways to search for multiple keywords.
The Elog search fields are treated as regular expressions.
In order to match a text containing "abc" and "def", you can use a search term "abc.*def".
A period (.) means "any character", and an asterisk (*) means "any number of repetition of the preceding character".
Therefore, ".*" matches "any number of any character" i.e. anything.
The search term "abc.*def" works fine when you know "abc" appears first in the text you are looking for.
If you don't know the order of appearance of the keywords, you have two choices: either to use,
"(abc.*def)|(def.*abc)"
or
"(abc|def).*(abc|def)"
The vertical bar (|) means "or". Parentheses are used for grouping.
The first example does exactly what you want. However, you have to list all the permutations of your keywords
separated by |. If you have more than two keywords, it can be a very very long search word.
(The length of the search word is O(n!), where n is the number of keywords).
In the second example, the length of the keyword is O(n). However, it can also match a text containing two "abc".
This means the search result may contain some garbages (entries containing only "abc").
I guess in most cases we can tolerate this.

To automatically construct a multiple keyword search term for the Elog, I wrote a bash script called elogkeywd
and it is installed in the control room machines.
You can type
elogkeywd keyword1 keyword2 keyword3
to generate a regular expression for searching a text containing "keyword1", "keyword2" and "keyword3".
The generated expression is of the second type shown above. You can then copy-and-paste the result to
the Elog search field.
The script takes any number of keywords. However, there seems to be a limit on the number of characters you can type
into the search field of the Elog. I found the practical limit is about 3 keywords.
  1180   Fri Dec 5 14:13:41 2008 ranaSummaryIOOMC trend for the last 4 days
The MC has stayed locked for ~3 days! I just broke it to reset the MZ.
Attachment 1: g.png
g.png
  1179   Fri Dec 5 09:29:59 2008 ranaUpdateIOOdrum modes observable without excitation
Not sure what the y-scale is since there aren't any y axis labels in the plot, but it seems like we
now get an SNR of a ~few with a BW of 0.1 Hz. IN principle, the frequency noise out of the PSL ought
to be limited by the VCO phase noise at these frequencies (sort of) so the broadband MC_F level
is very roughly equal to 20-100 mHz/rHz.

Since dnu = dL*(c/lambda)/L_MC, the thermal peaks have a height of ~10^-15 m_RMS. We (Caryn) should check
that these numbers are true and then see if this is the correct amount of energy for thermally excited
mirror modes.
  1178   Fri Dec 5 01:58:58 2008 YoichiConfigurationASCtdscntr.pl now works at 40m
Tobin gave me the perl version of tdscntr some time ago.
Pinkesh and I modified and tested it at LHO.
I further modified it today and now it runs fine on the linux machines at the 40m. I haven't tested it with the Solaris machines.
My modifications include changing channel names to 40m ones, and using tdsavg to get QPD data rather than ezcaread.
The use of tdsavg is intended to avoid aliasing problem.
tdscntr.pl is installed in /cvs/cds/caltech/apps/linux/tds/bin

Now, the alignX runs on linux up to the centering of the QPDs.
However, ezcademod seems to behave wrongly on linux. I plan to investigate on this problem tomorrow.
I may try tdsdmd instead.
  1177   Fri Dec 5 01:41:33 2008 YoichiConfigurationComputersMEDM screen snapshot now works on linux machines
As a part of my "make everything work on linux" project, I modified 'updatesnap' script so that linux machines can update MEDM screen snapshots.
Now, all 'updatesnap' in the subsystem directories (like medm/c1/lsc/cmd/updatesnap) are sym-link to /cvs/cds/caltech/medm/c1/cmd/updatesnap.
This script will take a window snapshot to a PNG file, and move the old snapshot to archive folders with date information added to the filename.
For compatibility, it also saves JPEG snapshot. Right now, most of 'view snapshot' menus in MEDM screens are calling 'sdtimage' command, which cannot display PNG files. I installed Imagemagick to op440m. We should change MEDM files to use 'display' command instead of 'sdtimage' so that it can show PNG files.
I've already changed some MEDM screens, but there are so many remaining to be modified.

PNG is better than JPEG for crisp images like screen shots. JPEG performs a sort of spacial Fourier transformations and low-pass filtering to compress the information. If it is used with sharp edges like boundaries of buttons on an MEDM screen, it naturally produces spacial aliasing (ghost images).

I also created several sym-links on the apps/linux/bin directory to mimic the Solaris-only commands, such as 'sdtimage', 'nedit' and 'dtterm'.
For example, nedit is symbolic linked to gedit. Many MEDM buttons/menus, which used to be incompatible with linux, now work fine on the linux machines.
  1176   Thu Dec 4 17:42:23 2008 carynUpdateIOOdrum modes observable without excitation
So, the mode cleaner was evidently aligned better and now the drum modes are observable using DTT.
The Lock-In was set to 27.8kHz and the drum mode frequencies were previously observed to be 28.039kHz(MC2), 28.222kHz(MC3) and 28.221kHz(MC1). So, we might expect peaks at ~239Hz,421Hz,422Hz.
Peaks have been observed around the expected frequencies in channel IOO-MC-DRUM1.
Note that it is possible to resolve the separate MC1 and MC3 peaks which are so close together.
(sorry these are pdf's and not png's)
Attachment 1: drum_modes.pdf
drum_modes.pdf
Attachment 2: drum_modes2.pdf
drum_modes2.pdf
  1175   Thu Dec 4 16:29:20 2008 josephbConfigurationComputersError message on Frame Builder Raid Array
The Fibrenetix FX-606-U4 RAID connected to the frame builder in 1Y7 is showing the following error message: IDE Channel #4 Error Reading
  1174   Thu Dec 4 13:49:39 2008 JenneUpdatePEMMore of: Comparing Wiener subtraction using different sensors
Here is another version of the same type plot I put in the elog yesterday. This plot is looking at the 7200 seconds after 04Dec2008 08:45:00 UTC. This time was last night, when there was no crazy seismic activity, and well after the Ranger seismometer was moved to its new place under MC2.

This plot includes all possible combinations of the accelerometers, Guralp seismometer and Ranger seismometer (taking all 6 accelerometers as a set, and all 3 Guralp channels as a set). It is good to see that for the set of traces which do not include the accelerometers - brown, dark green and light blue - the subtraction at higher frequencies isn't all that great. Thus, the accelerometers are doing their job, and work well with the Wiener subtraction.

Still under investigation is why we don't see a whole lot of improvement at low frequency.
Attachment 1: Dec042008_c1wino_seisCombos.png
Dec042008_c1wino_seisCombos.png
  1173   Wed Dec 3 20:36:07 2008 Jenne, RanaUpdatePEMComparing Wiener subtraction with different seismometers
The Ranger has now been moved over to sit underneath the MC2 tank (it was previously close to the PSL rack). It
is still pointed in the +Y direction (towards ETMY, aka south).

New spectra attached - looks like the coherence is still there between the Guralp and the Ranger which are now
seperated by the MC length (~12 m). At LLO, I have witnessed a coherence of less than 0.3 above 1 Hz for these
distances. Curious.

L960019-00-F describes measurements done at SLAC on seismic coherence. The iLIGO LSC PDD
(http://www.ligo.caltech.edu/docs/T/T970122-00.pdf) discusses in sec 4.2 how this was incorporated into the LSC design.

When we get our next Guralp, it will be interesting to move them around and determine what the cross-spectrum
is between different points in the lab during typical times.

In the second attachment, I have plotted the square of the quantity used in the LSC PDD (S_xy) which I think is
what we now plot in DTT as 'Coherence'.

The third attachment shows the coherences among the TM SUSPOS_INs. I've turned off the oplev servos for this but
the OSEM damping is still on. Its not quite the same as the theory, but we could probably measure/tweak the
seismic velocity and then get better agreement.
Attachment 1: d.pdf
d.pdf
Attachment 2: sco.png
sco.png
Attachment 3: fly.pdf
fly.pdf
  1172   Wed Dec 3 20:10:09 2008 Jenne, RanaUpdatePEMComparing Wiener subtraction with different seismometers
Attached is a plot of MC_L, and then the residual MC_L after static Wiener filtering, using different combinations of our accelerometers and seismometers.

This is the same type of plot that Rana has included in the past few weeks, using Wiener filters calculated with c1wino.m

This data is from GPS 912312914, duration = 7200 sec, sometime during the night last night.

Unfortunately, it doesn't look like adding the Guralp seismometer to the Accelerometers and the Ranger did much, especially at low frequencies (all sensors = black curve). We'll have to investigate why this is true, and what we can do to get some low-frequency subtraction going on.

In the legend, "Residuals Accels, Guralp, Ranger" implies that the residual has been calculated using all of the sensors listed.
Attachment 1: Dec032008_c1wino_seisCombos.png
Dec032008_c1wino_seisCombos.png
  1171   Wed Dec 3 19:21:09 2008 ranaConfigurationPEMRanger move
I looked at the Ranger signals. Somehow it has a relative transfer function of 'f' between it and the Guralp.
      Ranger
i.e.  ------ ~ f
      Guralp

which is strange since according to their manuals, they should both be giving us a voltage output which is proportional
to velocity. I checked that the Ranger only has a load resistor and then an SR560 low pass at 300 Hz. Jenne assures
me that the Guralp breakout box shouldn't have any poles either (to be double checked). Its a mystery.

We made sure that the SR560 now is DC coupled, G = 100, & 1-pole low pass at 300 Hz. I moved it over next to the Guralp
(went through the mass recentering procedure after forgetting to lock it before moving). It is behaving as it was
before.

Attached is a 2 page PDF of the comparisons. The 'MC1' channels are Guralp and 'MC2' is Ranger.

The second attachment compares our seismometers (in counts) with the LHO Guralp seismometers. There's no high frequency
rolloff there like what we see here so I bet a dollar that there's a pole in the Guralp box somewhere.
Attachment 1: c.pdf
c.pdf c.pdf
Attachment 2: wsnb.pdf
wsnb.pdf
  1170   Wed Dec 3 12:49:11 2008 jenneUpdateComputerssomething sketchy with NDS ... or something
Never mind...I had forgotten that you have to run mdv_config every time you open matlab, not just every time you boot a computer.

I am not able to get channels using get_data from the mDV toolbox on Allegra, Megatron or Rosalba.

The error I get while running the "hello_world" test program is:
hello_world
setting up configuration...
added paths for nds
added paths for qscan
couldn't add path for matapps_SDE
couldn't add path for matapps_path
couldn't add path for framecache
couldn't add path for ligotools_matlab
added paths for home_pwd
fetching channels for C...
Warning: get_channel_list() failed.
??? Error using ==> NDS_GetChannels
Failed to get channel list.

Error in ==> fetch_nds at 47
eval(['CONFIG.chl.' server ' = NDS_GetChannels(ab);']);

Error in ==> get_data at 100
out = fetch_nds(channels,dtype,start_time,duration);

Error in ==> hello_world at 6
aa = get_data('C1:LSC-DARM_ERR', 'raw', gps('now - 1 hour'), 32);
  1169   Wed Dec 3 11:58:10 2008 AlbertoUpdatePSLMC Alignment
Rana, Alberto,

more details on the MC alignment we did yesterday.


Last week Rana re-aligned the Mach Zender (MZ) on the PSL table to reduce the power at the dark port (see elog entry #1151). After that, the beam was aligned to the MZ but not properly aligned to the Mode Cleaner (MC) anymore. As a result the MC could not lock or did it on unwanted transverse modes. To fix that we decided to change the alignment of the MC input periscope on the PSL table.


The ultimate goal of the operation was to align the MC transmitted beam to the IFO and to maximize the power.
Such a condition depends on:
a) a good cavity alignment and
b) input beam matching to the cavity TEM00 mode.


Since the MZ alignment had only affected the input beam, we assumed the cavity alignment was still good, or at least it had not changed, and we focused on the input beam.

The IOO computer, by the MC autolocker script, is able to change the cavity alignment and the length to match the input beam and lock the cavity. Although both the length servo (LSC) and the alignment servo (WFS) have a limited effective operating range. So for the script to work properly and at best, input beam and cavity matching have to be not far from that range.

The MC periscope has two mirrors which control the pitch and yaw input angles. By changing either yaw or pitch of both mirrors together (“two-knob" technique) one can change the input angle without moving the injection point on the cavity input mirror (MC1). So this is the procedure that we followed:

  • 1) turned of the autolocker running the MC-down script
    2) brought the reflected beam spot back on the MC-reflection camera and on the reflection photodiode (REFL-PD)
    3) turned on the LSC servo
    4) tweaked the periscope's mirrors until the cavity got locked on a TEM00 mode
    5) tweaked the periscope aiming at ~0.3V from the REFL-PD and ~3V on the transmission photodiode (TRANS-PD).


Following the steps above we got ~0.5 V on the REFL-PD and ~2V on the TRANS-PD but no better than that.

Looking at the Drift Monitor MEDM screen, we found that the cavity was not in the reference optimal position, as we initially assumed, thus limiting the matching of the beam to the MC.

We restored the optics reference position and repeated the alignment procedure as above. This time we got ~3V on the TRANS-PD and ~0.5 on the REFL-PD. We thought that the reason for still such a relatively high reflection was that the beam was not well centered on the REFL-PD (high order modes pick-up?).

On the AS table we centered the REFL-PD by aligning a beam splitter in the optical path followed by the light to reach the photodiode.

We also centered the beam on the reflection Wave Front Sensors (WFS). To do that we halved the power on the MZ to reduce the sidebands power and prevent the WFS QPD from saturating. We then aligned the beam splitters on the QPD by balancing the power among the quadrants. Finally we restored the power on the MZ.

As a last thing, we also centered the transmitted beam on the TRANS-QPD.


The MC is now aligned and happily locked with 3V at the TRANS-PD and 0.3V at REFL-PD.
  1168   Tue Dec 2 19:51:32 2008 ranaUpdatePEMhalf-micron particle count is alarming
The 0.5 micron dust monitor count is now pretty high (36000). I wandered around the lab to see if there was anything
nasty going on but I didn't see or smell anything in particular. Since today Alberto was sitting around where the
dust monitor is while aligning the PSL beam, we should blame him. Its either garlic, cologne, or time to bathe.

The 400 day hour trend shows that while the counts are not so unusual, the 40m is dirtier than it was last year.
Attachment 1: Untitled.png
Untitled.png
Attachment 2: dust.png
dust.png
  1167   Tue Dec 2 19:18:10 2008 ranaSummaryPEMRanger SS-1
In entry http://dziban.ligo.caltech.edu:40/40m/881 and a follow up from Jenne I put in the Ranger calibration.
Since then, we've reduced the SR560 gain from 200 to 100 so the calibration factor is now:

1e-9 (m/s)/count and then 2 poles at 0 Hz, and a Q~1 zero pair at 1 Hz.
in DTT:
G = 1e-9
p = 0, 0
z = 0.7 0.7
  1166   Tue Dec 2 17:56:56 2008 Alberto, RanaConfigurationPSLMC Alignment
In the attempt to maximize the Mode Cleaner transmission and minimize the reflection from the steering mirrors of the MC periscope, we could not get more ~2 V at the MC Trans PD and ~ 0.5 V at MC REFL_DC. As it turned out from the SUS Drift Monitor, the reason was that the MC optics had been somehow displaced from the optimal position.

After restoring the reference position values for the mirrors and tweaking again the periscope, we got ~3V at the MC TransPD and 0.5V at the reflection.
The beam was then probably clipped at the REFL PD so that we had to adjust the alignment of one of the BS in the transmitted beam path on the AS table.
We also zeroed the WFS PDs, but not before reducing the power from the MZ, for their QPDs not to saturate.

After relocking, the transmission was 3V and the reflection ~0.3V.

The beam isnow centered on the Trans PD and REFL PD and the Mode Cleaner locked. More details on the procedure will follow.
  1165   Mon Dec 1 15:09:27 2008 robUpdatePEMhalf-micron particle count is alarming
  1164   Thu Nov 27 22:56:42 2008 ranaConfigurationEnvironmenttemperature
8-)
Attachment 1: mc.png
mc.png
  1163   Tue Nov 25 19:29:15 2008 rana, alberto, johnConfigurationEnvironmenttemperature

Quote:
The PSL Room Temperature was alarming because it had gone above 23 C. This set off an unfortunate chain of events:

We found that the PSL HEPA was set low (20%). This is a fine setting for when no one is working in there but it
does raise the temperature since there are heat sources inside the blue box.

We tried to change the office area temperature to compensate and also the westmost sensor inside the lab area by 2 deg F.

The office area one was problematic - there was so much dust in it that the gas valve nipple was clogged. So we've
now blown it all clean with a compressed air can. We're now tuning the calibration screw to make our new
digital sensor agree with the setpoint on the controller.

Expect wild temperature swings of the office area for a couple days while Alberto and I tune the servo.


This morning Bob found 92F in the office area and in the control room of the lab. He turned down the thermostat and when I got in at about 9 I found 65. After a few hours of adjustment of the thermostat's calibration, I could stabilize the room temperature to about 72F. I also turned down the thermostats inside the lab of a couple of degrees F.
  1162   Tue Nov 25 18:38:03 2008 Alberto, RobUpdatePSLMC Periscope Alignment
This morning when I came in I found the MC cleaner unlocked and the autolocker script could not lock it. The reflected beam was quite off and showed in the bottom left corner of the IMCR camera. After turning off the WFS locking, I started slightly changing the alignment of the steering mirrors on the MC periscope, waiting for the LSC servo to lock the cavity. It didn't work. At some point I lost the beam from the IMCR camera and that is how someone might have found it when I left it for about one hour.

When I came back and tried again adjusting the steering mirrors, I noticed that the autolocker was working and was trying to lock the cavity. After just a bit of adjustment, the MC got easily locked.

After that, I spent a couple of hours trying to improve the alignment of the periscope to minimize the reflection and maximize the transmission. I started with a transmission of 0.4 V but, despite all the tweaking (I used the technique of turning both yaw knobs at the same time), I couldn't get more than 1.2 V (and 2.4 V at the reflection) if only the LSC servo was on. Looking at the camera, I moved the beam around to look for a more favorable spot but the MC wouldn't lock with the beam in other places. Maybe I could do better or maybe not because the cavity is not aligned. I'm going to try again tomorrow.
  1161   Mon Nov 24 19:15:16 2008 rana, alberto, johnConfigurationEnvironmenttemperature
The PSL Room Temperature was alarming because it had gone above 23 C. This set off an unfortunate chain of events:

We found that the PSL HEPA was set low (20%). This is a fine setting for when no one is working in there but it
does raise the temperature since there are heat sources inside the blue box.

We tried to change the office area temperature to compensate and also the westmost sensor inside the lab area by 2 deg F.

The office area one was problematic - there was so much dust in it that the gas valve nipple was clogged. So we've
now blown it all clean with a compressed air can. We're now tuning the calibration screw to make our new
digital sensor agree with the setpoint on the controller.

Expect wild temperature swings of the office area for a couple days while Alberto and I tune the servo.
  1160   Mon Nov 24 17:14:44 2008 ranaUpdatePSLMach Zender trends
It looks like the MZ has gotten less drifty after the alignment on Friday.
Attachment 1: Untitled.png
Untitled.png
  1159   Mon Nov 24 16:43:34 2008 ranaConfigurationComputersAlex and Jay took away some computers from the racks
I was over at Wilson house and saw Jay and Alex bring in 3 rackmount computers. One was a Sun 4600 and
then there were 2 3U black boxes. I got the impression that these were the data concentrators or
data collectors or framebuilder test boxes. They said that they got these from the 40m and no one was
in the lab to oppose them except for Bob and he didn't put up much of a fight.

Everything looks green on the DAQ Detail and RFM network screens so perhaps everything is OK. Beware.
ELOG V3.1.3-